OpenStudy (anonymous):

1-(cos^(x))/(1+sin(x))

OpenStudy (anonymous):

i need help simplifying this. i have a photo to match the problem

OpenStudy (anonymous):

except the cos should be cos^2

OpenStudy (anonymous):

i thought i answered this. oops i was wrong sorry. numerator is \[sin^2(x)\]

OpenStudy (anonymous):

can you see the attachment?

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

\[{1-\cos^2x \over 1+\sin x}?\]

OpenStudy (anonymous):

no, its 1- the entire fraction i guess cos^2x/1+sinx

OpenStudy (anonymous):

ahhhhhhhh

OpenStudy (anonymous):

so numerator is \[1-sin^2(x)=(1+sin(x))(1-sin(x))\]

OpenStudy (anonymous):

cancel with the denominator to get \[1-sin(x)\]

OpenStudy (anonymous):

Oh I see. So it's \(1-\frac{\cos^2x}{1+sin x}\). \[=1-{1-\sin^2 x \over 1+\sin x}=1-{(1+\sin x)(1-\sin x) \over 1+\sin x}=1-1+\sin x=\sin x\]

OpenStudy (anonymous):

and \[1-(1-sin(x))=sin(x)\]

OpenStudy (anonymous):

what anwar said. hi answar!

OpenStudy (anonymous):

haha thank you for your help, sorry for the confusion

OpenStudy (anonymous):

Anwar :) .. and hi. Wow you have 517 medals satellite!